OpenStudy (anonymous):

Write the standard form of the equation of the line that passes through the point(4,1) and is parallel to the line 3x+3y=5

OpenStudy (anonymous):

3(x+y)=5 x+y=5/3 y=-x+5/3 so slope is -1.coefficient of x is alwas the slope now requird equn will be...(y-y1)=m(x-x1) where(x1,y1) is the given point and m is the slope

OpenStudy (anonymous):

maam/sir.. i guess i already got the answer

OpenStudy (anonymous):

can i ask one more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Solve the given linear system 2x-y-z=5 x+y+z=7 3x-2y-3z=1

OpenStudy (anonymous):

u need it by which method?

OpenStudy (anonymous):

uhm.. substitution..?

OpenStudy (anonymous):

2x-5=y+z now put the value of y+z in seocnd equation, x+(2x-5)=7

OpenStudy (anonymous):

can u solve dis one and find the value of x?

OpenStudy (anonymous):

i already found the value of x

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

but the thing is i dont know what to do next..

OpenStudy (anonymous):

see,now put the value of x in any two equations and solve for y and z.first substitute in any 2 eqns

OpenStudy (anonymous):

uhmm can you do it for me and explain how?

OpenStudy (anonymous):

i m tconsidering these 2 equations, x+y+z=7 first we put value of x in dis eqn, 4+y+z=7 y+z=7-3 y+z=4.........(1) another eqn is, 3x-2y-3z=1 now putting the value of x in dis eqn,we get 3*4-2y-3z=1 12-1=2y+3z 11=2y+3z.........(2)

OpenStudy (anonymous):

now u jst need to solve these 2 eqns and get the value of y and z

OpenStudy (anonymous):

how did the 4 became 3 in eq. 1?

OpenStudy (anonymous):

i still can't get it :/

OpenStudy (anonymous):

can you solve it?

OpenStudy (anonymous):

for me? so i can answer some of the items for tomorrows long quiz? :/

OpenStudy (anonymous):

multiply eqn 1 by 2,we get 2y+2z=8 now subtracting dis eqn frm 2nd eqn,

OpenStudy (anonymous):